SJSU’s student help center has been monitoring number of students who come in to seek support with their basic English course. Below you’ll find the number of students that have gone to the student help center during the last eight weeks.

1

2

3

4

5

6

7

8

20

26

31

33

39

42

48

52

Produce a 3-week moving average forecast for the Periods 4–8.

Group of answer choices

29.5, 32, 26, 40.5, 44

29.4, 34, 38, 42.4, 42

25.7, 30, 34.3, 38, 48

28.5, 32, 36, 40.5, 45
Answer #1

Forecast = Sigma(Demand in previous N periods) / N
Where N = 3

FORECAST 4 = (20 + 26 + 31) / 3 = 25.7
FORECAST 5 = (26 + 31 + 33) / 3 = 30
FORECAST 6 = (31 + 33 + 39) / 3 = 34.3
FORECAST 7 = (33 + 39 + 42) / 3 = 38
FORECAST 8 = (39 + 42 + 48) / 3 = 43